# Predictive Control and Regret Analysis of Non-Stationary MDP with Look-ahead Information

### v1

@@ 6.2 EV charging
In this section, we consider an EV charging station under the setup of Example 2. The charging station has three charging stands, and the energy price fluctuates between 2 and 18.
We first show the correlation between the change of energy price and energy usage. We compare the regret of our algorithm with the benchmark policy smoothed least-laxity-first algorithm (sLLF) proposed in Chen et al. [2022], which prorize charging the EV that is the closest to the departure time. However, given the fluctuating energy price, the optimal policy should charge the EVs at the time steps with the lowest energy price that can still satisfy the energy demand of the EVs before their departure times. As shown in Figure 3(a), compared with sLLF, our algorithm selects better time for charging each EV. In particular, when $k$ increases, most of the peak of energy demand falls within the shaded area with low energy price.
We then show the decay in regret with respect to the growth in the prediction horizon. Compared with traditional scheduling algorithm proposed in Chen et al. [2022], our algorithm can better handle the fluctuation in energy price. As shown in Figure 3(b), even with only a few steps of prediction, the station’s regret decays exponentially.
(a) Power usage at different time steps. The shaded area indicates the time period with energy price below 8. We see that with with $k\geq 7$, the most energy usage happens within the area with low energy price, reducing the total energy cost of the station.
(b) The above plot shows that the regret of EV charging decays with the prediction horizon. Compared with traditional scheduling policies, the proposed algorithm can lower the total energy cost even with a few prediction steps.
Figure 3:
@@ Appendix A Contraction in span semi-norm
In this section, we briefly introduce some properties of span semi-norm, and how it helps to show the contraction property in Theorem 4.1.
@@
The proof of above proposition easily follows from 2. More detailed properties of the span semi-norm can be found in Puterman [1994]. In the following proposition, we show the fundamental step of $J$-stage contraction.
@@
Let $v\in\mathbb{R}^{n}$ and $\pi_{t},\dots,\pi_{t+J}\in\mathcal{A}^{J}$, then
@@
$=\frac{1}{2}\max_{s,u\in\mathcal{S}}\sum_{j\in\mathcal{S}}|P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|s)-P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|u)|=\max_{s,u\in\mathcal{S}}\sum_{j\in\mathcal{S}}[P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|s)-P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|u)]^{+}$
@@
We are now ready to prove Proposition 3.1.
Proof:  [Proof of Proposition 3.1] Let $v_{t}^{1},v_{t}^{2}\in\mathbb{R}^{|\mathcal{S}|}$ denote any pair of arbitrary vectors. $s^{*}=\arg\max_{s\in\mathcal{S}}\{L_{t:t+J}v_{t}^{1}(s)-L_{t:t+J}v_{t}^{2}(s)\}$, and $s_{*}=\arg\min_{s\in\mathcal{S}}\{L_{t:t+J}v_{t}^{1}(s)-L_{t:t+J}v_{t}^{2}(s)\}$. Further, let $\pi_{t+i}^{1}$ denote the optimal policy according to $L_{t+i}\circ\cdots\circ L_{t+J}(v_{t}^{1})$, and let $\pi_{t+i}^{2}$ denote the optimal policy to take according to $L_{t+i}\circ\cdots\circ L_{t+J}(v_{t}^{2})$ then
$L_{t:t+J}v_{t}^{1}(s^{*})-L_{t:t+J}v_{t}^{2}(s^{*})$ $\leq L_{t:t+J}^{\pi_{t:t+J}^{1}}v_{t}^{1}(s^{*})-L_{t:t+J}^{\pi_{t:t+J}^{1}}v_{t}^{2}(s^{*})=P_{t:t+J}^{\pi_{t:t+J}^{1}}(v_{t}^{1}-v_{t}^{2})(s^{*})$
$L_{t:t+J}v_{t}^{1}(s_{*})-L_{t:t+J}v_{t}^{2}(s_{*})$ $\geq L_{t:t+J}^{\pi_{t:t+J}^{2}}v_{t}^{1}(s_{*})-L_{t:t+J}^{\pi_{t:t+J}^{2}}v_{t}^{2}(s_{*})=P_{t:t+J}^{\pi_{t:t+J}^{2}}(v_{t}^{1}-v_{t}^{2})(s_{*})$
@@
$\operatorname{sp}(L_{t:t+J}v_{t}^{1}-L_{t:t+J}v_{t}^{2})$ $\leq P_{t:t+J}^{\pi_{t:t+J}^{1}}(v_{t}^{1}-v_{t}^{2})(s^{*})-P_{t:t+J}^{\pi_{t:t+J}^{2}}(v_{t}^{1}-v_{t}^{2})(s_{*})$
$\leq\max_{s\in\mathcal{S}}P_{t:t+J}^{\pi_{t:t+J}^{1}}(v_{t}^{1}-v_{t}^{2})(s)-\min_{s\in\mathcal{S}}P_{t:t+J}^{\pi_{t:t+J}^{2}}(v_{t}^{1}-v_{t}^{2})(s)$
$\leq\operatorname{sp}\left(\begin{bmatrix}P_{t:t+J}^{\pi_{t:t+J}^{1}}\\ P_{t:t+J}^{\pi_{t:t+J}^{2}}\end{bmatrix}(v_{t}^{1}-v_{t}^{2})\right)$ (16)
Applying Proposition A.2 to (16) immediately leads to the theorem statement. $\square$
@@ Appendix B Diameter
In this section, we prove Proposition 3.3, which shows that the span semi-norm of value function $V_{t}$ is upper bounded by diameter $D$ for all $t$.
Proof:  [Proof of Proposition 3.3] Let $\pi^{*}$ denote the optimal policy and $\pi^{\prime}$ denote the policy defined in Definition 4 trying to move fastest to $\{s_{t+i}^{*}\}_{i\in[D]}$. By Definition 4, under a trajectory generated by $\pi^{\prime}$ starting from time $t$ at $s_{*}$, define $d=\inf\{\tau:\tau>0,s_{t+\tau}=s_{t+\tau}^{*}\}$. $d$ is a stopping time and by Definition 4, $\mathbb{E}[d]\leq D$.
@@
$=$ $\mathbb{E}^{\pi^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s^{*}]-\mathbb{E}^{\pi^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s_{*}]$
$\leq$ $\mathbb{E}^{\pi^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s^{*}]-\mathbb{E}^{\pi^{\prime}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s_{*}]$
$=$ $\left(\mathbb{E}^{\pi^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})|s_{t}=s^{*}]-\mathbb{E}^{\pi^{\prime}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})|s_{t}=s_{*}]\right)$
$+\underbrace{\left(\mathbb{E}^{\pi^{*}}[V_{t+d}^{*}(s_{t+d})|s_{t}=s^{*}]-V_{t+d}^{*}(s_{t+d}^{*})\right)}_{<0}$
$\leq$ $\mathbb{E}d\leq D$
@@
In this section, we prove Lemma 5.2, which bounds the step-wise error incurred by using the forecast system dynamics, instead of the ground-truth transition and reward functions.
@@
The first inequality is due to the relative optimality of $\hat{a}$ for $\hat{V}$ , and the second/third inequalities are by 1 and Holder’s inequality. Similarly,
@@
Combining (17) and (18), we obtain
$\begin{split}&r_{t+\ell}(s,\tilde{a})+\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\tilde{V}(s^{\prime})-\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)-\epsilon_{\ell}-\delta_{\ell}D\\ \leq&\hat{r}_{t+\ell}(s,\hat{a})+\mathbb{E}_{s^{\prime}\sim\hat{P}_{t+\ell|t}(\cdot|s,\hat{a})}\hat{V}(s^{\prime})\\ \leq&r_{t+\ell}(s,\tilde{a})+\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\tilde{V}(s^{\prime})-\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\hat{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)+\epsilon_{\ell}+\delta_{\ell}D\end{split}$ (19)
@@
Substituting (20) into (19), we obtain
$-\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)-\epsilon_{\ell}-\delta_{\ell}D\leq\hat{L}_{t+\ell|t}\hat{V}(s)-L_{t+\ell}\tilde{V}(s)\leq-\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\hat{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)+\epsilon_{\ell}+\delta_{\ell}D.$ (21)
Let $s^{*}:=\arg\max_{s}\{\hat{L}_{t+\ell|t}\hat{V}(s)-L_{t+\ell}\tilde{V}(s)\}$ and $s_{*}:=\arg\min_{s}\{\hat{L}_{t+\ell|t}\hat{V}(s)-L_{t+\ell}\tilde{V}(s)\}$. Using 2 and the assumption that $\operatorname{sp}\left(\tilde{V}-\hat{V}\right)<b$, we obtain
@@ Appendix D Bounding the error of $Q$ function
$\operatorname{sp}\left(\hat{\psi}_{t}^{0}-\tilde{\psi}_{t}^{0}\right)\leq 2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right)$
@@
$\tilde{\psi}_{t}^{0}-\hat{\psi}_{t}^{0}=L_{t}\circ\dots\circ L_{t+k}W_{0}-\hat{L}_{t|t}\circ\dots\circ\hat{L}_{t+k|t}W_{0}$ (23)
We proceed step-wise on $\operatorname{sp}(\hat{\psi}_{t}^{\ell}-\tilde{\psi}_{t}^{\ell})$ for $\ell$ going backward from $k$ to 00. For the base case where $\ell=k+1$, we have $\operatorname{sp}\left(\tilde{\psi}_{t}^{k+1}-\hat{\psi}_{t}^{k+1}\right)=\operatorname{sp}(W_{0}-W_{0})=0$.
@@
$\leq 2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right)$ (26)
where (24) is obtained by Lemma 5.2, and (25) is obtained by Proposition 3.1. We repeat the steps in (24) and (25) to obtain (26). $\square$
@@
If $t+k\geq T$, $\tilde{\psi}_{t}^{k}=V_{t}^{*}$.
Proof:   The latter equality is clear from the definition of Bellman operator, so we just need to prove the first inequality.
@@
$=\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+1}^{*}\right)$
where we applied Proposition 3.1 for $\lfloor k/J\rfloor$ times. The last equality holds because $W_{0}=0$. $\square$
@@
For any states $s$ and $t<T-k$, the $Q$ function of the MDP satisfies
@@
Proof:   By the Bellman operator,
@@
$\leq$ $\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+1}^{*}\right)$
where we used Lemma D.2 for the last inequality. $\square$
Then, we bound the difference between the $\tilde{\Psi}_{t}$ and the estimated $Q$ function $\hat{\Psi}_{t}$ with forecast under 1.
@@
$|\tilde{\Psi}_{t}(s,a)-\hat{\Psi}_{t}(s,a)|<\epsilon_{0}+\delta_{0}D+2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right)$
Proof:
$|\tilde{\Psi}_{t}(s,a)-\hat{\Psi}_{t}(s,a)|\leq$ $|r_{t}(s,a)-\hat{r}_{t|t}(s,a)|+\left|\mathbb{E}_{s^{\prime}\sim P_{t}(\cdot|s,a)}[\tilde{\psi}_{t+1}^{0}(s^{\prime})]-\mathbb{E}_{s^{\prime}\sim\hat{P}_{t}(\cdot|s,a)}[\hat{\psi}_{t+1}^{0}(s^{\prime})]\right|$
$\leq$ $\epsilon_{0}+\left|\mathbb{E}_{s^{\prime}\sim P_{t}(\cdot|s,a)}[\tilde{\psi}_{t+1}^{0}(s^{\prime})-\hat{\psi}_{t+1}^{0}(s^{\prime})]\right|+\delta_{0}D$

### v2

@@ 6.2 EV charging
Figure 4: Figure˜4(a) shows the power usage at different time steps. The shaded area indicates the time period with energy prices below 8. We see that with $k\geq 7$, the most energy usage happens within the area with low energy prices, reducing the total energy cost of the station. Figure˜4(b) demonstrates that the regret of EV charging decays with the prediction horizon. Compared with traditional scheduling policies, the proposed algorithm can lower the total energy cost even with a few prediction steps.
In this section, we consider a scenario of EV charging station under the setup of Example˜2 with time horizon $T=50$. The charging station has three charging stands, and the energy price fluctuates between 2 and 18.
We first show the correlation between the change of energy price and energy usage. We compare the regret of our algorithm with the benchmark policy, the smoothed least-laxity-first algorithm (sLLF) proposed in Chen22, which prioritizes charging the EV that is the closest to the departure time. However, given the fluctuating energy price, the optimal policy should charge the EVs at the time steps with the lowest energy price that can still satisfy the energy demand of the EVs before their departure times. As shown in Figure˜4(a), compared with sLLF, our algorithm selects better time for charging each EV. In particular, when $k$ increases, most of the peak of energy demand falls within the shaded area with low energy price.
We then show the decay of regret with respect to the growth in the prediction horizon. Compared with traditional scheduling algorithm proposed in Chen22, our algorithm can better handle the fluctuation in energy price. As shown in Figure˜4(b), even with only a few steps of prediction, the station’s regret decays exponentially.
@@ Appendix A Contraction in span semi-norm
In this section, we briefly introduce some properties of span semi-norm, and how it helps to show the contraction property in Theorem˜4.1.
@@
The proof of the above proposition easily follows from ˜1. More detailed properties of the span semi-norm can be found in Puterman94. In the following proposition, we show the fundamental step of $J$-stage contraction.
@@
Let $v\in\mathbb{R}^{n}$ and policies $\pi_{t},\dots,\pi_{t+J}$, then
@@
$=\frac{1}{2}\max_{s,u\in\mathcal{S}}\sum_{j\in\mathcal{S}}|P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|s)-P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|u)|=\max_{s,u\in\mathcal{S}}\sum_{j\in\mathcal{S}}[P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|s)-P_{t:t+J}^{\pi_{t}:\pi_{t+J}}(j|u)]^{+}.$
@@
We are now ready to prove Proposition˜3.1.
Proof: [Proof of Proposition˜3.1] Let $v^{*}$ denote the optimal value function at time $t+J+1$, and let $v\in\mathbb{R}^{|\mathcal{S}|}$ denote the value function for an arbitrary policy at time $t+J+1$. $s^{*}=\arg\max_{s\in\mathcal{S}}\{L_{t:t+J}v^{*}(s)-L_{t:t+J}v(s)\}$, and $s_{*}=\arg\min_{s\in\mathcal{S}}\{L_{t:t+J}v^{*}(s)-L_{t:t+J}v(s)\}$. Further, let $\pi_{t+i}^{*}$ denote the optimal policy according to $L_{t+i}\circ\cdots\circ L_{t+J}(v^{*})$, and let $\pi_{t+i}$ denote the optimal policy to take according to $L_{t+i}\circ\cdots\circ L_{t+J}(v)$ then
$L_{t:t+J}v^{*}(s^{*})-L_{t:t+J}v(s^{*})$ $\leq L_{t:t+J}^{\pi_{t:t+J}^{*}}v^{*}(s^{*})-L_{t:t+J}^{\pi_{t:t+J}^{*}}v(s^{*})=P_{t:t+J}^{\pi_{t:t+J}^{*}}(v^{*}-v)(s^{*}),$
$L_{t:t+J}v^{*}(s_{*})-L_{t:t+J}v(s_{*})$ $\geq L_{t:t+J}^{\pi_{t:t+J}^{2}}v^{*}(s_{*})-L_{t:t+J}^{\pi_{t:t+J}}v(s_{*})=P_{t:t+J}^{\pi_{t:t+J}}(v^{*}-v)(s_{*}).$
@@
$\operatorname{sp}(L_{t:t+J}v^{*}-L_{t:t+J}v)$ $\leq P_{t:t+J}^{\pi_{t:t+J}^{*}}(v^{*}-v)(s^{*})-P_{t:t+J}^{\pi_{t:t+J}}(v^{*}-v)(s_{*})$
$\leq\max_{s\in\mathcal{S}}P_{t:t+J}^{\pi_{t:t+J}^{*}}(v^{*}-v)(s)-\min_{s\in\mathcal{S}}P_{t:t+J}^{\pi_{t:t+J}}(v^{*}-v)(s)$
$\leq\operatorname{sp}\left(\begin{bmatrix}P_{t:t+J}^{\pi_{t:t+J}^{*}}\\ P_{t:t+J}^{\pi_{t:t+J}}\end{bmatrix}(v^{*}-v)\right).$ (16)
Applying Proposition A.2 to equation 16 immediately leads to the theorem statement. $\square$
@@ Appendix B Diameter
In this section, we prove Proposition˜3.3, which shows that the span semi-norm of value function $V_{t}$ is upper bounded by diameter $D$ for all $t$.
Proof: [Proof of Proposition˜3.3] Let $\bm{\pi}^{*}$ denote the optimal policy and $\bm{\pi}^{\prime}$ denote the policy defined in Definition 3 trying to move fastest to $\{s_{t+i}^{*}\}_{i\in[D]}$. By Definition 3, under a trajectory generated by $\bm{\pi}^{\prime}$ starting from time $t$ at $s_{*}$, define $\inf\{\tau:\tau>0,s_{t+\tau}=s_{t+\tau}^{*}\}. dd$ is a stopping time and by Definition 3, $\mathbb{E}[d]\leq D$.
@@
$=$ $\mathbb{E}^{\bm{\pi}^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s^{*}]-\mathbb{E}^{\bm{\pi}^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s_{*}]$
$\leq$ $\mathbb{E}^{\bm{\pi}^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s^{*}]-\mathbb{E}^{\bm{\pi}^{\prime}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})+V_{t+d}^{*}(s_{t+d})|s_{t}=s_{*}]$
$=$ $\left(\mathbb{E}^{\bm{\pi}^{*}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})|s_{t}=s^{*}]-\mathbb{E}^{\bm{\pi}^{\prime}}[\sum_{h=t}^{t+d-1}r_{h}(s_{h},a_{h})|s_{t}=s_{*}]\right)$
$+\underbrace{\left(\mathbb{E}^{\bm{\pi}^{*}}[V_{t+d}^{*}(s_{t+d})|s_{t}=s^{*}]-V_{t+d}^{*}(s_{t+d}^{*})\right)}_{<0}$
$\leq$ $\mathbb{E}d\leq D.$
@@
In this section, we prove Lemma˜5.2, which bounds the step-wise error incurred by using the forecast system dynamics, instead of the ground-truth transition and reward functions.
@@
The first inequality is due to the relative optimality of $\hat{a}$ for $\hat{V}$ , and the second/third inequalities are by ˜1 and Holder’s inequality. Similarly,
@@
Combining equation 17 and equation 18, we obtain
$\begin{split}&r_{t+\ell}(s,\tilde{a})+\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\tilde{V}(s^{\prime})-\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)-\epsilon_{\ell}-\delta_{\ell}D\\ \leq&\hat{r}_{t+\ell}(s,\hat{a})+\mathbb{E}_{s^{\prime}\sim\hat{P}_{t+\ell|t}(\cdot|s,\hat{a})}\hat{V}(s^{\prime})\\ \leq&r_{t+\ell}(s,\tilde{a})+\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\tilde{V}(s^{\prime})-\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\hat{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)+\epsilon_{\ell}+\delta_{\ell}D.\end{split}$ (19)
@@
Substituting equation 20 into equation 19, we obtain
$\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\tilde{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)-\epsilon_{\ell}-\delta_{\ell}D\leq\hat{L}_{t+\ell|t}\hat{V}(s)-L_{t+\ell}\tilde{V}(s)\leq-\mathbb{E}_{s^{\prime}\sim P_{t+\ell}(\cdot|s,\hat{a})}\left(\tilde{V}(s^{\prime})-\hat{V}(s^{\prime})\right)+\epsilon_{\ell}+\delta_{\ell}D.$ (21)
Let $s^{*}:=\arg\max_{s}\{\hat{L}_{t+\ell|t}\hat{V}(s)-L_{t+\ell}\tilde{V}(s)\}$ and $s_{*}:=\arg\min_{s}\{\hat{L}_{t+\ell|t}\hat{V}(s)-L_{t+\ell}\tilde{V}(s)\}$. Using ˜1 and the assumption that $\operatorname{sp}\left(\tilde{V}-\hat{V}\right)<b$, we obtain
@@ Appendix D Bounding the error of $Q$ function
$\operatorname{sp}\left(\hat{\psi}_{t}^{0}-\tilde{\psi}_{t}^{0}\right)\leq 2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right),$
@@
$\tilde{\psi}_{t}^{0}-\hat{\psi}_{t}^{0}=L_{t}\circ\dots\circ L_{t+k}W_{0}-\hat{L}_{t|t}\circ\dots\circ\hat{L}_{t+k|t}W_{0}.$ (23)
We proceed step-wise on $\operatorname{sp}(\hat{\psi}_{t}^{\ell}-\tilde{\psi}_{t}^{\ell})$ for $\ell$ going backward from $k$ to 0. For the base case where $\ell=k+1$, we have $\operatorname{sp}\left(\tilde{\psi}_{t}^{k+1}-\hat{\psi}_{t}^{k+1}\right)=\operatorname{sp}(W_{0}-W_{0})=0$.
@@
$\leq 2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right),$ (26)
where equation 24 is obtained by Lemma˜5.2, and equation 25 is obtained by Proposition˜3.1. We repeat the steps in equation 24 and equation 25 to obtain equation 26. $\square$
@@
If $t+k\geq T$, the estimation of $V_{t}^{*}$ is exact, i.e. $\tilde{\psi}_{t}^{k}=V_{t}^{*}$.
Proof: The latter equality is clear from the definition of the Bellman operator, so we just need to prove the first inequality.
@@
$=\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+1}^{*}\right),$
where we applied Proposition˜3.1 for $\lfloor k/J\rfloor$ times. The last equality holds because $W_{0}=0$. $\square$
@@
For any states $s$ and $t+k<T$, the $Q$ function of the MDP satisfies
@@
If $t+k\geq T$, the estimation of $Q_{t}^{*}$ is exact, i.e. $\tilde{\Psi}_{t}(s,\cdot)=Q_{t}^{*}(s,\cdot)$.
Proof: The latter equality is a direct result of the construction of $\tilde{\Psi}_{t}$ in equation 14, so we only need to prove the first inequality. By the Bellman operator,
@@
$\leq$ $\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+1}^{*}\right),$
where we used Lemma˜D.2 for the last inequality. $\square$
Then, we bound the difference between the $\tilde{\Psi}_{t}$ and the estimated $Q$ function $\hat{\Psi}_{t}$ with forecast under ˜1.
@@
$\operatorname{sp}\big(\tilde{\Psi}_{t}(s,\cdot)-\hat{\Psi}_{t}(s,\cdot)\big)<2\epsilon_{0}+2\delta_{0}D+2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right).$
Proof: By triangle inequality in Proposition˜A.1, we have
$\operatorname{sp}\big(\tilde{\Psi}_{t}(s,\cdot)$ $-\hat{\Psi}_{t}(s,\cdot)\big)\leq\operatorname{sp}\big(r_{t}(s,\cdot)-\hat{r}_{t|t}(s,\cdot)\big)+\operatorname{sp}\Big(\mathbb{E}_{s^{\prime}\sim P_{t}(\cdot|s,\cdot)}[\tilde{\psi}_{t+1}^{0}(s^{\prime})]-\mathbb{E}_{s^{\prime}\sim\hat{P}_{t}(\cdot|s,\cdot)}[\hat{\psi}_{t+1}^{0}(s^{\prime})]\Big)$
$\leq$ $2\epsilon_{0}+\operatorname{sp}\Big(\mathbb{E}_{s^{\prime}\sim P_{t}(\cdot|s,\cdot)}[\tilde{\psi}_{t+1}^{0}(s^{\prime})-\hat{\psi}_{t+1}^{0}(s^{\prime})]\Big)+\operatorname{sp}\Big(\mathbb{E}_{s^{\prime}\sim P_{t}(\cdot|s,\cdot)}[\hat{\psi}_{t+1}^{0}(s^{\prime})]-\mathbb{E}_{s^{\prime}\sim\hat{P}_{t}(\cdot|s,\cdot)}[\hat{\psi}_{t+1}^{0}(s^{\prime})]\Big)$
$\leq$ $2\epsilon_{0}+\operatorname{sp}\Big(\mathbb{E}_{s^{\prime}\sim P_{t}(\cdot|s,\cdot)}[\tilde{\psi}_{t+1}^{0}(s^{\prime})-\hat{\psi}_{t+1}^{0}(s^{\prime})]\Big)+2\delta_{0}D$